OpenStudy (anonymous):

A 1560-kg car is traveling with a speed of 16.4 m/s. What is the magnitude of the horizontal net force that is required to bring the car to a halt in a distance of 52.1 m?

OpenStudy (anonymous):

using v^2=u^2+2as you get a then using F=ma you get the required F

OpenStudy (anonymous):

ok thanks